# Reflection Theory of Nichols Algebras over Coquasi-Hopf Algebras with Bijective Antipode

### v1

@@ 4.1 Hopf pairings for Nichols algebras
Let $\mathcal{C}$ be an abelian braided monoidal category. We first recall the definition of Nichols algebras $\mathcal{B}(V)$ in $\mathcal{C}$, where $V\in\mathcal{C}$ is an arbitrary object. Then we are going to show that there is a non-degenerate Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$ when $V$ has a left dual $V^{*}$.
@@
where $n]!_{c}\in\operatorname{End}(V^{\otimes n})$ is the braided symmertriser. For explicit definition of $n]!_{c}$, one may refer to [Bazlov and Berenstein, 2013, Section 5.6].
@@
We are going to investigate the Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$ when $V$ has a left dual $V^{*}$. There is a unique Hopf pairing $\omega:T(V^{*})\otimes T(V)\to\textbf{1}$. It is shown in [Majid, 1993] that $\omega$ is given by
@@
$\omega_{n}:=\omega|_{V^{*\otimes n}\otimes V^{\otimes n}}=\operatorname{ev}_{V^{\otimes n}}\circ(\operatorname{id}_{V^{*\otimes n}}\otimes[n]!_{c})=\operatorname{ev}_{V^{\otimes n}}\circ([n]!_{c^{*}}\otimes\operatorname{id}_{V^{\otimes n}})$
The Hopf pairing between $V^{\otimes n}$ and $V^{*\otimes m}$ is 0 unless $n=m$. Here $c^{*}$ denotes the the braided symmertriser of $V^{*}$.
The Hopf pairing between $T(V)$ and $T(V^{*})$ inducing a Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$. Note that the Hopf pairing $\omega$ vanishes on the kernels of the symmetrisers, and for all positive integer $n$, we have
@@
Now we return to the case of coquasi-Hopf algebra $H$ with bijective antipode. Let $V$ be a finite-dimensional object in $^{H}_{H}\mathcal{YD}$}. Consequently, the Hopf pairing on the tensor algebras descends to a well-defined Hopf pairing
@@
Let $A$, $B$ be $\mathbb{N}_{0}$-graded locally finite Hopf algebra in $\mathcal{C}$ with bijective antipodes and $\omega:A\otimes B\rightarrow\mathbbm{k}$ a dual pair. Recall the braided monoidal equivalence in Theorem 3.16:
@@
If $V$ is rational, then for any $n\in\mathbb{N}_{\geq 1}$, $V^{\otimes n}$ is rational in $^{B}_{B}\mathcal{YD}(\mathcal{C})$}. Since ${{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$ is closed under direct sum and quotient by Lemma 3.3 both objects $T(V)=\bigoplus_{n\geq 0}V^{\otimes n}$ and $\mathcal{B}(V)$ are rational. ∎
@@ 4.2 The functor $\Omega$ under $\mathbb{Z}$-gradings
[…]
Furthermore, suppose $H$ is a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra. A $\mathbb{Z}$-graded Yetter-Drinfeld module $V$ over $H$ is by definition an object $V$ in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$. In other words, $V$ is a $\mathbb{Z}$-graded vector space such that the map $H\otimes V\rightarrow V$ and comodule structure maps $V\rightarrow H\otimes V$ are $\mathbb{Z}$-graded.
@@
Now we return to the purpose of this subsection. We first introduce the following notations for further use.
@@
$\lambda_{\mathcal{B}(N^{*})}=(\mathcal{B}(N^{*})\otimes V\xrightarrow{\operatorname{id}\otimes\delta_{\mathcal{B}(N)}}\mathcal{B}(N^{*})\otimes(\mathcal{B}(N)\otimes V)\xrightarrow{a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}}(\mathcal{B}(N^{*})\otimes\mathcal{B}(N))\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}\mathcal{B}(N^{*})),$
@@
is $\bigoplus_{m\neq n}\mathcal{B}(N^{*})(m)\otimes V$. This means, if $v\in\mathcal{F}^{n}V$, then for each $i>n$, $\lambda_{\mathcal{B}(N^{*})}(\mathcal{B}(N^{*})(i)\otimes v)=0$ by (4.3). On the other hand, if $v\in\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$. which completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})=\mathcal{F}^{n}V$.
@@
5 Reflection of Nichols algebras over coquasi-Hopf algebra with bijective antipode
Let $H$ be a coquasi-Hopf algebra with bijective antipode. In this section, we study projection of Nichols algebras in $^{H}_{H}\mathcal{YD}$}, which will play an important role when considering reflection of simple Yetter-Drinfeld modules. Some proofs in this section are parallel to those in the pointed coquasi-Hopf algebra setting Li and Liu (2025b), we will omit them for simplicity.
@@ 5.1 The structure of the space of coinvariants $K$
Denote $\pi_{H}:\mathcal{A}(M\oplus N)\rightarrow H$, and $\pi^{\prime}_{H}:\mathcal{A}(N)\rightarrow H$ be the natural projection. These maps satisfy a compatibility condition, namely, there exists a commutative diagram relating these projections:
@@
Here, $\mathbb{S}$ denotes preantipode of $\mathcal{A}(N)$. We have the following two observations.
@@
The element $\operatorname{ad}(a)(x)$ is of degree $\operatorname{deg}(X)+1$ in the Nichols algebra $\mathcal{B}(M\oplus N)$. Moreover, we have $\operatorname{ad}(N)^{n}(M)\cap\operatorname{ad}(N)^{m}(M)=0$ for $n\neq m$. This is implies the direct sum decomposition of $L$.
@@
Now we assume that for some fixed integer $n_{0}$, we have $\delta_{K}(\operatorname{ad}(a)(x))\in\mathcal{A}(N)\otimes L$ for all $x\in\operatorname{ad}(N)^{n^{\prime}}(M)$, where $n^{\prime}\leq n_{0}$. Then for $a\in N$ and $x\in\operatorname{ad}(N)^{n_{0}+1}(M)$. Since by definition, $\operatorname{ad}(N)^{n_{0}+1}(M)=\operatorname{ad}(N)(\operatorname{ad}(N)^{n_{0}}(M))$. We have $\delta_{K}(x)\in\mathcal{A}(N)\otimes L$ by inductive hypothesis. Thus $\operatorname{ad}(a_{0})(x_{0})\in\operatorname{ad}(N)(L)\subseteq L$. On the other hand, $\operatorname{ad}(a_{-2})(x_{0})\in L$ as well. Since $\operatorname{ad}(a_{-2})(x_{0})\in\bigoplus_{i=0}^{n_{0}+1}\operatorname{ad}(N)^{i}(M)$, we have $(\operatorname{ad}(a_{-2})(x_{0}))_{0}\in L$ by inductive hypothesis. This completes the induction, and we shows $\delta_{K}:L\rightarrow\mathcal{A}(N)\otimes L$ is well-defined. The third axiom holds automatically since $L\subseteq K$. Thus $L$ is a $\mathbb{N}_{0}$-graded object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} and is a subobject of $K$.
@@
Since the term $x\otimes 1$ lies in $\Delta_{\mathcal{B}(M\oplus N)}(x)$, let us consider this expression separately. The term $\operatorname{ad}(a)(x)\otimes 1$ will occur in $\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(a)(x))$. Now the left terms are
@@
Now we combine the second and third term together, yields
@@
(1) It follows direct that $K$ inherits an $\mathbb{N}_{0}$-graded Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $K(n)=K\cap(\mathcal{B}(M\oplus N)\#H)(n)$.
@@
(3) Note that $K(1)=K\cap(\mathcal{B}(M\oplus N)\#H)(1)=K\cap(\mathcal{B}(M\oplus N)(1))=\operatorname{ad}(\mathcal{B}(N))(M)=L$. We are going to show $K$ is generated its degree one component $L$. Let $K^{\prime}$ be the subalgebra of $K$ generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, which is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with structures induced by tensor product of $L$. Now we let $W$ be the image of $K^{\prime}\#\mathcal{B}(N)$ under the isomorphism $K\#\mathcal{B}(N)\cong\mathcal{B}(M\oplus N)$. It suffices to show that $W=\mathcal{B}(M\oplus N)$. Note that $M\oplus N\subseteq W$ since $M,N\subseteq W$. Thus, we need only verify that $W$ forms a subalgebra of $\mathcal{B}(M\oplus N)$.
@@
Since $\operatorname{ad}(a)(x)\in K^{\prime}$ and $\operatorname{ad}(a_{-1})(x)a_{0}\in K^{\prime}\#\mathcal{B}(N)$, we conclude that $ax\in K^{\prime}\#\mathcal{B}(N)$. Hence its image lies in $W$. Thus $W$ is a subalgebra of $\mathcal{B}(M\oplus N)$, which establishes that $K$ is generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
@@
${{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. By [Li and Liu, 2025b, Lemma 4.12], $M$ is an object in $^{H}_{H}\mathcal{YD}$} with decomposition into irreducible objects $M=\bigoplus_{i\in I}M_{i}:=\bigoplus_{i\in I}M\cap W_{i}$. Moreover, we have $W_{i}=\mathcal{B}(N)\rhd M_{i}=\bigoplus_{n\in\mathbb{N}_{0}}N^{\otimes n}\rhd M_{i}$ for each $i\in I$.
Let $\operatorname{deg}(M_{i})=m_{i}$ for each $i\in I$. We endow $W_{i}$ a new grading by $\widetilde{W_{i}}=W_{i}$ with
@@
is a $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Its degree 0’s part is $\mathbbm{k}$, and its degree $1$’s part is $M\oplus N$.
@@
Since $M\in{{}^{H}_{H}\mathcal{YD}}$ and $y_{-1}\in H$, the term $y\rhd x$ is contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Since $\mathcal{B}(N)\rhd M$ is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, $\delta(y\rhd x)\in\mathcal{A}(N)\otimes(N\rhd M)$. This implies terms like $(y\rhd x)_{0}$ are contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Let us fixed $k\geq 0$, for all $z\in\mathcal{B}(N)(k^{\prime})$ and $x\in M$, where $k^{\prime}\geq k$. We assume the element $z\rhd x$ as well as $(z\rhd x)_{0}$ lie in the subalgebra generated by $\mathcal{B}(N)$ and $M$, then for homogeneous $y\in N$,

### v2

@@ 4.1 Hopf pairings for Nichols algebras
Let $\mathcal{C}$ be an abelian braided monoidal category. We first recall the definition of Nichols algebras $\mathcal{B}(V)$ in $\mathcal{C}$, where $V\in\mathcal{C}$ is an arbitrary object. Then we will show that there is a non-degenerate Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$ when $V$ has a left dual $V^{*}$.
@@
where $n]!_{c}\in\operatorname{End}(V^{\otimes n})$ is the braided symmertriser. For the explicit definition of $n]!_{c}$, one may refer to [Bazlov and Berenstein, 2013, Section 5.6].
@@
We are going to investigate the Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$ when $V$ has a left dual $V^{*}$. There is a unique Hopf pairing $\omega:T(V^{*})\otimes T(V)\to\textbf{1}$. It is shown in [Bazlov and Berenstein, 2013] that $\omega$ is given by
@@
$\omega_{n}:=\omega|_{V^{*\otimes n}\otimes V^{\otimes n}}=\operatorname{ev}_{V^{\otimes n}}\circ(\operatorname{id}_{V^{*\otimes n}}\otimes[n]!_{c})=\operatorname{ev}_{V^{\otimes n}}\circ([n]!_{c^{*}}\otimes\operatorname{id}_{V^{\otimes n}}).$
The Hopf pairing between $V^{\otimes n}$ and $V^{*\otimes m}$ is 0 unless $n=m$. Here $c^{*}$ denotes the braided symmetriser of $V^{*}$.
The Hopf pairing between $T(V)$ and $T(V^{*})$ induces a Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$. Note that the Hopf pairing $\omega$ vanishes on the kernels of the symmetrisers, and for all positive integer $n$, we have
@@
Now we return to the case of a coquasi-Hopf algebra $H$ with bijective antipode. Let $V$ be a finite-dimensional object in $^{H}_{H}\mathcal{YD}$}. Consequently, the Hopf pairing on the tensor algebras descends to a well-defined Hopf pairing
@@
Let $A$, $B$ be $\mathbb{N}_{0}$-graded locally finite Hopf algebra in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ with bijective antipodes and $\omega:A\otimes B\rightarrow\mathbbm{k}$ a dual pair. Recall the braided monoidal equivalence in Theorem 3.16:
@@
If $V$ is rational, then for any $n\in\mathbb{N}_{\geq 1}$, $V^{\otimes n}$ is rational in $^{B}_{B}\mathcal{YD}(\mathcal{C})$}. Since ${{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$ is closed under direct sum and quotient by Lemma 3.3. Both objects $T(V)=\bigoplus_{n\geq 0}V^{\otimes n}$ and $\mathcal{B}(V)$ are rational. ∎
@@ 4.2 The functor $\Omega$ under $\mathbb{Z}$-gradings
Having established the duality for Nichols algebras, we now investigate how the functor $\Omega$ interacts with gradings. Let $\mathbb{N}_{0}\text{-Gr}\mathcal{M}_{\mathbbm{k}}$(resp., -$\mathbb{Z}\text{-Gr}\mathcal{M}_{\mathbbm{k}}$) be the category of $\mathbb{N}_{0}$-graded vector spaces(resp., $\mathbb{Z}$-graded vector spaces). The functor $\mathbb{N}_{0}\text{-Gr}\mathcal{M}_{k}\rightarrow\mathbb{Z}\text{-Gr}\mathcal{M}_{k}$, which extends the $\mathbb{N}_{0}$-grading of an object $V$ in $\mathbb{N}_{0}\text{-Gr}\mathcal{M}_{k}$ to a $\mathbb{Z}$-grading by setting $V(n)=0$ for all $n<0$, is monoidal. This functor allows us to view $\mathbb{N}_{0}$-graded coalgebras and coquasi-Hopf algebras as $\mathbb{Z}$-graded coalgebras and $\mathbb{Z}$-graded coquasi-Hopf algebras, respectively.
Furthermore, suppose $H$ is a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra. A $\mathbb{Z}$-graded Yetter-Drinfeld module $V$ over $H$ is by definition an object $V$ in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$. In other words, $V$ is a $\mathbb{Z}$-graded vector space such that the map $H\otimes V\rightarrow V$ and the comodule structure maps $V\rightarrow H\otimes V$ are $\mathbb{Z}$-graded.
@@
Now we return to the purpose of this subsection. We first introduce the following notation for further use.
@@
$\lambda_{\mathcal{B}(N^{*})}=(\mathcal{B}(N^{*})\otimes V\xrightarrow{\operatorname{id}\otimes\delta_{\mathcal{B}(N)}}\mathcal{B}(N^{*})\otimes(\mathcal{B}(N)\otimes V)\xrightarrow{a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}}(\mathcal{B}(N^{*})\otimes\mathcal{B}(N))\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}V),$
@@
is $\bigoplus_{m\neq n}\mathcal{B}(N^{*})(m)\otimes V$. This means, if $v\in\mathcal{F}^{n}V$, then for each $i>n$, $\lambda_{\mathcal{B}(N^{*})}(\mathcal{B}(N^{*})(i)\otimes v)=0$ by (4.3). On the other hand, if $v\in\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$. This completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})=\mathcal{F}^{n}V$.
@@
5 Reflection of Nichols algebras over coquasi-Hopf algebras with bijective antipode
Let $H$ be a coquasi-Hopf algebra with bijective antipode. In this section, we study projection of Nichols algebras in $^{H}_{H}\mathcal{YD}$}, which will play an important role when considering reflection of simple Yetter-Drinfeld modules. Some proofs in this section are parallel to those in the pointed coquasi-Hopf algebra setting Li and Liu (2025b); therefore we will omit them for simplicity.
@@ 5.1 The structure of the space of coinvariants $K$
Let $\pi_{H}:\mathcal{A}(M\oplus N)\rightarrow H$, and $\pi^{\prime}_{H}:\mathcal{A}(N)\rightarrow H$ be the natural projection. These maps satisfy a compatibility condition, namely, there exists a commutative diagram relating these projections:
@@
Here, $\mathbb{S}$ denotes the preantipode of $\mathcal{A}(N)$. We have the following two observations.
@@
The element $\operatorname{ad}(a)(x)$ is of degree $\operatorname{deg}(X)+1$ in the Nichols algebra $\mathcal{B}(M\oplus N)$. Moreover, we have $\operatorname{ad}(N)^{n}(M)\cap\operatorname{ad}(N)^{m}(M)=0$ for $n\neq m$. This implies the direct sum decomposition of $L$.
@@
Now we assume that for some fixed integer $n_{0}$, we have $\delta_{K}(\operatorname{ad}(a)(x))\in\mathcal{A}(N)\otimes L$ for all $x\in\operatorname{ad}(N)^{n^{\prime}}(M)$, where $n^{\prime}\leq n_{0}$. Then for $a\in N$ and $x\in\operatorname{ad}(N)^{n_{0}+1}(M)$, since by definition $\operatorname{ad}(N)^{n_{0}+1}(M)=\operatorname{ad}(N)(\operatorname{ad}(N)^{n_{0}}(M))$, we have $\delta_{K}(x)\in\mathcal{A}(N)\otimes L$ by the inductive hypothesis. Thus $\operatorname{ad}(a_{0})(x_{0})\in\operatorname{ad}(N)(L)\subseteq L$. On the other hand, $\operatorname{ad}(a_{-2})(x_{0})\in L$ as well. Since $\operatorname{ad}(a_{-2})(x_{0})\in\bigoplus_{i=0}^{n_{0}+1}\operatorname{ad}(N)^{i}(M)$, we have $(\operatorname{ad}(a_{-2})(x_{0}))_{0}\in L$ by inductive hypothesis. This completes the induction, and we show $\delta_{K}:L\rightarrow\mathcal{A}(N)\otimes L$ is well-defined. The third axiom holds automatically since $L\subseteq K$. Thus $L$ is a $\mathbb{N}_{0}$-graded object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} and is a subobject of $K$.
@@
Since the term $x\otimes 1$ lies in $\Delta_{\mathcal{B}(M\oplus N)}(x)$, let us consider this expression separately. The term $\operatorname{ad}(a)(x)\otimes 1$ will occur in $\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(a)(x))$. Now the remaining terms are
@@
Combining the second and third terms together yields
@@
(1) It follows directly that $K$ inherits an $\mathbb{N}_{0}$-graded Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $K(n)=K\cap(\mathcal{B}(M\oplus N)\#H)(n)$.
@@
(3) We are going to show $K$ is generated by $L$. Let $K^{\prime}$ be the subalgebra of $K$ generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, which is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with structures induced by tensor product of $L$. Now we let $W$ be the image of $K^{\prime}\#\mathcal{B}(N)$ under the isomorphism $K\#\mathcal{B}(N)\cong\mathcal{B}(M\oplus N)$. It suffices to show that $W=\mathcal{B}(M\oplus N)$. Note that $M\oplus N\subseteq W$ since $M,N\subseteq W$. Thus, we need only verify that $W$ forms a subalgebra of $\mathcal{B}(M\oplus N)$.
@@
Since $\operatorname{ad}(a)(x)\in K^{\prime}$ and $\operatorname{ad}(a_{-1})(x)a_{0}\in K^{\prime}\#\mathcal{B}(N)$, we conclude that $ax\in K^{\prime}\#\mathcal{B}(N)$. Hence its image lies in $W$. Thus $W$ is a subalgebra of $\mathcal{B}(M\oplus N)$, which establishes that $K$ is generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. Note that $L\subseteq K(1)$, this forces $K(1)=L$. Hence $K$ is generated by $K(1)$.
@@
${{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. By [Li and Liu, 2025b, Lemma 4.12], $M$ is an object in $^{H}_{H}\mathcal{YD}$} with decomposition into irreducible objects $M=\bigoplus_{i\in I}M_{i}:=\bigoplus_{i\in I}M\cap W_{i}$. Here $M_{i}$ was proven to be the $\mathbb{Z}$-homogeneous component of $W_{i}$ of minimal degree. Moreover, we have $W_{i}=\mathcal{B}(N)\rhd M_{i}=\bigoplus_{n\in\mathbb{N}_{0}}N^{\otimes n}\rhd M_{i}$ for each $i\in I$.
Denote $\operatorname{deg}(M_{i})=m_{i}$ for each $i\in I$. We endow $W_{i}$ with a new grading by $\widetilde{W_{i}}=W_{i}$ with
@@
is an $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Its degree 0 part is $\mathbbm{k}$, and its degree $1$ part is $M\oplus N$.
@@
Since $M\in{{}^{H}_{H}\mathcal{YD}}$ and $y_{-1}\in H$, the term $y\rhd x$ is contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Since $\mathcal{B}(N)\rhd M$ is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, $\delta(y\rhd x)\in\mathcal{A}(N)\otimes(N\rhd M)$. This implies terms like $(y\rhd x)_{0}$ are contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Let us fix $k\geq 0$, for all $z\in\mathcal{B}(N)(k^{\prime})$ and $x\in M$, where $k^{\prime}\geq k$. We assume the element $z\rhd x$ as well as $(z\rhd x)_{0}$ lie in the subalgebra generated by $\mathcal{B}(N)$ and $M$, then for homogeneous $y\in N$,